Tech Staff Meeting Notes September 19, 2008

Current Top Priorities are: Station maintenance


There are four levels of priority.
PrioritiesWish to try to get these done (or make progress on) before the next meeting (yeah right)
Second TierFit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month
As Time PermitsFit these in when you can but not instead of first or second priorities
OtherStuff we hope to get to and don't want to forget about


The task list updated February 6, 2008. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.


*************************************************


Priorities

Station Maintenance, NMSZ

Station maintenance is the top priority.

Several of the stations can get improved telemetry by transmitting directly to the node. The original configuration changed enough so that there is room in the band. Jim will investigate.
Holly's list of problem stations:

Isis Oxidation problem
BRNM.EHZ
SJBM.EHN
NHAR.EHE
JHAR.EHE
PPLM.EHZ
MCAM.EHN

Down or unusable
PLAR.EHN
PLAR.EHE
QUAR.EHE
TMAR.EHN
LRAR.EHE
CPAR.EHZ
TNMT.EHN
MIST.EHE
CHRM
DWDM.EHE
MLDM

There was extensive storm damage from Ike in the southeast MO region including several seismic stations. Will begin the repairs as soon possible.

TIPT replacement. NWCT is schedule for completion this week. ETN

CCRT - Needs a rebuild after a bad lightning strike.

NC Node - Need to figure out how to fix GPS time drop-outs. Jim suggests making sure the clock is configured correctly and not alarming on inconsequential hiccups.

GFM - Fixed the bad channel but still needs some work to repair storm damage.

RBNC - Replace battery(s), and add another solar panel.

GMG - Need battery box solar reg mounting plate and need to fuse the battery.

Need to acquire and install new batteries for all the XLAK clocks.

Routine Maintenance Kudzu, battery check, etc, etc.


*************************************************


Second Tier Priorities

Replace batteries in all XLAK clocks. Need to be careful so that earthworm doesn't see time jump around.

Most of the strongmotion stations can run at 100 sps without any modifications. We will install a router on the few that have problems (RDST will be the test case). We will also migrate the acquisition software from gcf2ew on anssworm to scream/scream2ew on a linux or ms server. The latter will facilitate stations maintenance without interrupting the data stream.

All CMG5TD sensors need to have the latest firmware (59914) installed and trigger download verified. Also need to adjust the trigger level as warranted by the local noise. The nominal level is 7000 "micro-g" which is really counts and translates to about 1% g (nominal conversion is 130,000 count/m/s/s). Note that some DAS's are fussy and you may need to set them to filing mode and then to dual to convince them to save triggers. So far these are done: NHIN, NMAD, LNXT, CVTN, and SFTN.

Need a new bottom air conditioner for mkta. Need new RF preamps, radio align, etc, etc at MKTA.
PEBM and HENM sometimes have telemetry fades. Should reevaluate comms.
Need to knock down the jungle around the guy anchors at NMAD.
Need to finish the tower work at NMAD.
Need to swap out the old node computers at NHIN, and AZTN. Will XP do netbui with the bridge recorders (if so we could lose the windows 98 box at aztn)?

GPS

ET

LRVA - East/West channel has low amplitude and needs troubleshooting.

WSNC - (Winespring Bald, NC) - Have Special use permit from Forest Service. Need to determine details of exactly what goes precisely where. John will get some pics and discuss via phone with Jim and Chris.

Fort Payne - John found a good spot with decent odds of a permission at Desoto state park in Alabama. It will have to repeat through a Georgia fire tower and John is investigating the permit there.

The map shows stations with yellow triangles being new ones.
stationman-daysstatus
MGNC8done
CCNC8done
WSSR8new
TVNC8done
LRVA8done
Maryville8new
VHTN8done
RBNC4done
TRYN4done
GFM4done
SMNC4done (see note above)
PLVA2abandon
BHT8move (eventually)

HDBR
The recorder disks should be reformatted as a routine maintenance task next time access is available. The constant read/write to the fat file systems lead to fragmentation. Don't ya just love dos?
Lost contact with recorders 1, 3, and 6. Number 5 is unstable. Can reset all 7 recorders which implies the equipment enclosure at each node has power and that at least the reset radio and telemetry are working.

HDAR
This is the point where we hop on the TDOT wireless network. Its also a strong motion station. When the funds come through, this will become an 18-component "station" like Mud Island. The surface station needs to move to where the boreholes will be. Chris met with AR DOT people and figured out where everything is supposed to go. TDOT will likely be ready for us in September.

HDBT
Mud Island vertical array is operational. Not really, functional is a better word until we can get it online to qc/vet data. Trigger levels on the strongmo need to be adjusted because the flash fills in about two weeks.

CVTN optical isolation. Need to upgrade to year 3 hardware. Hardware ready.

Steve needs to speak with Dean Kemp at NWS. They are willing to let us piggyback on their cable internet (not .nws.noaa.gov).

Lenox Agreement. Jim obtained contact info for the person Mitch needs to call>

FDL. This is the FDL Policy. .
Horton and Langston deployed the CMG6TD's to record Gustav and found that 3 of them were not operational. Steve and Dave will get together to go over all of them to "check-in" the ones that are ok and to assess what the others need to make them operational.
Horton also reports that "There are six GMT6TD cables that need attention. Four are firewire cables that require a water tight cap to cover the exposed end. The water seal on two other connectors need to be inspected and repaired if needed." David is looking into it.

For the record, Jim gave this end of summer report on August 27:
Here's update from the end of the summer and last couple of weeks. -New Madrid node:
HICK-possible lightning damage repaired
BROM-ISIS canister rebuilt and reinstalled
CATM-antenna problem repaired
All other stations serviced (minus PARM no access and 3 ISIS in TN) and many deferred maintenance issues fixed.
Priorities-NCCT install, SJBM canister rebuild, central receive tower work and BACM
-Lenox node:
GLAT-DM24 hung?, now back up
GOBM-solar regulator repaired
All other stations serviced and many deferred maintenance issues fixed.
Several days of extensive generator maintenance done (propane tank filled?)
Priorities-Several ISIS single component outages, hook generator status lines into lnxtworm and rework CWPT telemetry
Marked Tree node:
All but 2 stations serviced and many deferred maintenance issues fixed.
Priorities-QUAR and CPAR component outages, ISIS telemetry hardware upgrade at CR & return spare AC units to both doghouses.


*************************************************


As Time Permits

Had Cummings in to repair new generator at 3904 central. Chris serviced the old generator.

Need to gather up a contact list (e.g. everyone's home phone, power companies, etc) and put it on webworm (password protected).

Can we route Pat's alarms into the earthworm digitizer? If they are just a flat -5/0/+5 voltage that monitors a state change, it should be easy to add them into the env_alarm channels.

CPCT - Re-work lighting and pest chaser. John has a pest chaser and will use some landscape or pool hardware for the lighting.

SWET - Need to seal the hinges on the aluminum Hoffman boxes with flashing tape to keep the water out.

3876 Power The power for the new lighting on the east side of the general parking lot to the west of CERI is coming from 3876. Chris will check it out to assess the added lightning hazard and see if anything can be done about it. Might be hard to do without the wiring diagram. Chris will make one if needed as time permits.

3890 Power Chris is going to make a diagram for 3890. Good luck son.

Inventory Mitch needs to look at this and determine where we are and what needs to be done to make this operational.


*************************************************


Other

Withers needs to determine current orientation for MPHB so we can pull it and orient correctly.

Pink Palace is working with Gary and Michelle on a new perc-like display area. We'll pull the receiver etc, when they start pulling out the old display (or sometime earlier if an opportunity arises).

Jim will switch the FCC licenses from Greg's name to Mitch.